Consider the reaction of 2 mol H_2(g) at 25 degree
Answer #1

2H2 + O2 ---> 2H2O

change in moles delta(n) = 2 - 3 = -1

T = 25 oC = 25 + 273 = 298 K

R = 8.314 J/K.mol

Feed values,

maximum work done = -delta(n)RT = -(-1) x 8.314 x 298 = 2.48 kJ

Change in entropy deltaS = -nRlnX

X being the mole fraction of H2 or O2

mole fraction of H2 = 2/3 = 0.67

mole fraction of O2 = 1/3 = 0.33

deltaS = -(2 x 8.314)ln0.67 - (1 x 8.314)ln0.33

           = 6.66 + 9.22

          = 15.88 J/K